Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2018-05-10

Observers:Owen McBrienPeter Clark
Support Team:David O'NeillKen Smith
TAT:Regis CartierJoe Lyman
Daytime Calibrations:
Flats, arcs, biasGR 11, 13, 16
Twilight Start:22:01
Evening Calibrations:
grzFinal z band exposure failedWeather:Hazy at sunset,
Twilight End:23:23
Morning Calibrations:
ArcsGR 11, 13, 16 Erratic seeing particualy at the start of the night
UTObjectObservation TypeSetupSlit WidthExp. time(s)AirmassDIMMImage SeeingCommentsClassification
22:44Image Analysis/Focus
23:04SN2018behFollowupAbortedTelescope suffered a guiding software fault-
23:49AT2018bfqClassificationGr#131''1x 9001.1-1.2Ia
00:47L745aStandardGr#11, Gr#13, Gr#161''800 total1.6-1.4-
01:15AT2018bdgClassiciationGr#131''1x 6001.8-No nearby starsSlit rotated slightly from paragalactic angle due to lack of guide starsIa
01:36SN2017hpiFollowupGr#161''2700Start 1.6, End1.72.2-
02:32AT2018bgbClassificationGr#131.5''1x 9001.251.6-1.21.2Ia
03:03SN2017ijnFollowupGr#111''1x18001.91.161.2-
03:49AT2018bhbClassiciationGr#131''1x9001.81.4 - 1.61.4Seeing worsened significantly during exposureIa
04:09SN2017cbvFollowupGr#131.5''1x 27001.11.4-2.01.6-
05:04SN2018ahqFollowupGr#131''1x 27001,41.31.3-
06:02EG274StandardGr#131'' + 1.5''1801.021.221.25-
06:20AT2018bhcClassificationGr#131''1x 6001.731.31.3Ia
06:37AT2018bcmClassificationGr#131''1x 15001.851.71.3Ia
07:09SN2017bzcFollowupGr#131''3x 27002.2,1.7,1.351.35,1.22,1.32-


Observations End:09:57
Twilight Start:09:55
Comments